OpenStudy (anonymous):

Solve: negative 1 over 2 x + 2 = -x + 7

sam (.sam.):

\[-\frac{1}{2x+2}=-x+7 ~~??\]

sam (.sam.):

-1=(-x+7)(2x+2) -1=-2x^2-2x+14x+14 -1=-2x^2+12x+14 2x^2-12x-15=0

OpenStudy (anonymous):

thanks for yur help but i dont think thats right

sam (.sam.):

I don't think your question is right, can you draw it out?